# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v2

@@ IV-A O-P Strategy
The traditional training and testing approach for CD models is illustrated in the upper-left corner of Fig. 3 (green box) and the upper-right corner (pink box). For a given CD model $\mathcal{M}$, the input consists of a pair of images $(X_{1},X_{2})$, and the output is a change map (CM). For binary CD tasks with a single channel $c=1$, the predicted label $\hat{Y}$ is typically determined using a thresholding method:
$\hat{Y}(i)=\begin{cases}1,&\text{if }CM(i)>th\\ 0,&\text{if }CM(i)\leq th\end{cases}$ (1)
where $i$ denotes the pixel location in the image, “1” represents a change, and “0” represents no change (for visualization purpose, “1” will be mapped to a grayscale value of 255). The threshold $th$ is a predefined value. If $c=2$, the predicted label $\hat{Y}$ is generally determined via an element-wise comparison between the two layers of the change map, as follows:
$\hat{Y}(i)=\begin{cases}1,&\text{if }CM_{0}(i)\geq CM_{1}(i)\\ 0,&\text{if }CM_{0}(i)<CM_{1}(i)\end{cases}$ (2)
where $CM_{0}$ and $CM_{1}$ represent the first and second layers of the change map, respectively. If $c>2$, the CD model $\mathcal{M}$ not only predicts the locations of changes but also detects the types of changes, which is beyond the scope of this paper.
However, as shown in Fig. 4, for a dataset like JL1-CD, where the Change Area Ratio (CAR) can range from 0% to 100%, training a single model using traditional methods may not be optimal. The model would struggle to learn the full range of change patterns in an all-inclusive manner. To address this issue, we propose the Origin-Partition (O-P) strategy to enhance the model’s detection performance. As illustrated in the green boxes of Fig. 3 (Step 1 and Step 2), for a given CD algorithm, we train the corresponding models in the following sequence:
@@
2) As shown in Fig. 5, the CAR range for the training, validation, and test sets is very large. However, the majority of images have a CAR of less than 5%. Therefore, we set appropriate thresholds $th_{1}$ and $th_{2}$ and divide the original training set into three categories: small, medium, and large.
@@
$\hat{Y}=\begin{cases}f_{\mathcal{M}_{T_{S}}}(X_{1},X_{2}),&\text{if }CAR_{GT}\leq th_{1}\\ f_{\mathcal{M}_{T_{M}}}(X_{1},X_{2}),&\text{if }th_{1}<CAR_{GT}\leq th_{2}\\ f_{\mathcal{M}_{T_{L}}}(X_{1},X_{2}),&\text{if }CAR_{GT}>th_{2}\end{cases}$ (3)
@@
where $CAR_{GT}$ denotes the CAR calculated based on the ground truth label for the image pair $(X_{1},X_{2})$.
@@
As shown in the middle pink box in Fig. 3, during testing, since we do not know the CAR of the test images, we first use $\mathcal{M}_{O}$ to estimate the CAR roughly. Based on this estimated CAR, we then classify the image into one of the three categories: small, medium, or large, and send it to the corresponding model $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, or $\mathcal{M}_{T_{L}}$ to obtain the final detection result:
$\hat{Y}=\begin{cases}f_{\mathcal{M}_{T_{S}}}(X_{1},X_{2}),&\text{if }CAR_{\mathcal{M}_{O}}\leq th_{1}\\ f_{\mathcal{M}_{T_{M}}}(X_{1},X_{2}),&\text{if }th_{1}<CAR_{\mathcal{M}_{O}}\leq th_{2}\\ f_{\mathcal{M}_{T_{L}}}(X_{1},X_{2}),&\text{if }CAR_{\mathcal{M}_{O}}>th_{2}\end{cases}$ (4)
@@
where $CAR_{\mathcal{M}_{O}}$ denotes the CAR calculated based on the predicted label from the original model $\mathcal{M}_{O}$.
IV-B MTKD Framework
@@
By partitioning the training set, we effectively reduce the learning burden for each model. As a result, the Origin-Partition (O-P) strategy significantly enhances the performance of the CD algorithm. However, the O-P strategy still has two significant limitations: First, during inference, we are required to load four different models, and even disregarding data throughput, the time required is at least twice that of the original algorithm (often much more), significantly increasing both memory and time complexity. Second, since we first use $\mathcal{M}_{O}$ to obtain a rough estimate of the CAR, any inaccuracies in this estimation may lead to incorrect model selection in the subsequent steps, thereby reducing the accuracy of the final predictions. We are thus prompted to consider: is there a way to combine the capabilities of these four models into a single model?
@@
To address this, we propose the Multi-Teacher Knowledge Distillation (MTKD) framework. In the O-P strategy, we have already trained the models $\mathcal{M}_{O}$, $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$. Building on this, we further train a student model $\mathcal{M}_{S}$. First, we initialize the student model $\mathcal{M}_{S}$ using the parameters from $\mathcal{M}_{O}$, and then use $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$ as teacher models to perform KD. For each input image pair, we select the appropriate teacher model based on the image’s CAR to guide the student model. The process is shown in the green box at the bottom of Fig. 3. In this framework, the student model $\mathcal{M}_{S}$ is simultaneously supervised by the ground truth labels and the CM information from the teacher models across different CAR partitions.
@@
During the testing phase, only the student model $\mathcal{M}_{S}$ is used for inference, thereby significantly improving the model’s CD performance across different CAR ranges without introducing any additional computational cost.
IV-C Loss Function
@@
When training the original model $\mathcal{M}_{O}$ and the teacher models $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$, we employ the standard binary cross-entropy loss:
$L_{\text{CE}}=\frac{1}{N}\sum_{i=1}^{N}(-Y(i)\log(\hat{Y}(i))-(1-Y(i))\log(1-\hat{Y}(i)))$ (5)
@@
When training the student model $\mathcal{M}_{S}$, we select the appropriate teacher model based on the CAR range. Then the mean squared error (MSE) is computed at the CM layer as the distillation loss:
$L_{\text{KD}}=\frac{1}{N}\sum_{i=1}^{N}\left(CM(i)-CM_{\mathcal{T}}(i)\right)^{2},\quad\mathcal{T}\in\{T_{S},T_{M},T_{L}\}.$ (6)
Thus, the complete loss function for training $\mathcal{M}_{S}$ is given by:
$L=L_{\text{CE}}+\lambda L_{\text{KD}}$ (7)
@@
where the parameter $\lambda$ is used to balance the contributions of the cross-entropy loss and the distillation loss.
TABLE II: Benchmark Methods and the Corresponding Implementation Details

### v3

@@ 4.1 O-P Strategy
The traditional training and testing paradigm for CD models is depicted in the upper-left (training, green box) and upper-right (testing, pink box) sections of Figure 3. In binary CD tasks, given a model $\mathcal{M}$ and an image pair $(X_{1},X_{2})$, the model outputs a change map (CM), from which the predicted label $\hat{Y}$ is derived via thresholding or element-wise comparison.
To quantify the diversity of changes within a dataset, we introduce the concept of Change Area Ratio (CAR), defined as the ratio of changed pixels relative to the total image area. A detailed discussion of the CAR distribution on the JL1-CD and SYSU-CD datasets is provided in Appendix A. For datasets like them, where CAR values span a wide range (0% to 100%), traditional single-stage training tends to yield models that struggle to perform well across all CAR scenarios. To address this issue, we propose a general Origin-Partition (O-P) strategy. Rather than training a single model on the entire training set, we partition the dataset into multiple subsets according to different CAR thresholds, which enables each model to focus on a specific range of change scenarios. As a concrete instantiation (see Figure  3, green boxes at Step 1 and Step 2), for a given CD algorithm, we train the corresponding models in the following sequence:
@@
2) To reduce the training complexity, we set appropriate CAR thresholds $th_{1}$ and $th_{2}$ to partition the original training set into three categories: small, medium, and large.
@@
$\hat{Y}=\begin{cases}f_{\mathcal{M}_{T_{S}}}(X_{1},X_{2}),&\text{if }\text{CAR}_{\text{GT}}\leq th_{1};\\ f_{\mathcal{M}_{T_{M}}}(X_{1},X_{2}),&\text{if }th_{1}<\text{CAR}_{\text{GT}}\leq th_{2};\\ f_{\mathcal{M}_{T_{L}}}(X_{1},X_{2}),&\text{if }\text{CAR}_{\text{GT}}>th_{2},\end{cases}$ (1)
@@
where $\text{CAR}_{\text{GT}}$ denotes the CAR calculated based on the ground truth label for the image pair $(X_{1},X_{2})$.
@@
As shown in the middle pink box in Figure 3, during testing, since the true CAR of the test images is unknown, we first use $\mathcal{M}_{O}$ to estimate the CAR roughly. Based on this estimated CAR, we then assign the images to one of the three categories: small, medium, or large, and send them to the corresponding model $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, or $\mathcal{M}_{T_{L}}$ to obtain the final detection result:
$\hat{Y}=\begin{cases}f_{\mathcal{M}_{T_{S}}}(X_{1},X_{2}),&\text{if }\text{CAR}_{\mathcal{M}_{O}}\leq th_{1};\\ f_{\mathcal{M}_{T_{M}}}(X_{1},X_{2}),&\text{if }th_{1}<\text{CAR}_{\mathcal{M}_{O}}\leq th_{2};\\ f_{\mathcal{M}_{T_{L}}}(X_{1},X_{2}),&\text{if }\text{CAR}_{\mathcal{M}_{O}}>th_{2},\end{cases}$ (2)
@@
where $\text{CAR}_{\mathcal{M}_{O}}$ denotes the CAR calculated based on the predicted label from the original model $\mathcal{M}_{O}$.
4.2 MTKD Framework
@@
Partitioning the training set in the O-P strategy effectively reduces the training complexity for each model and can enhance CD performance to some extent. Nevertheless, this approach suffers from two major limitations. First, during inference, we are required to load four different models, and even disregarding data throughput, the time required is at least twice that of the original method, significantly increasing both memory and computational complexity. Second, since we first use $\mathcal{M}_{O}$ to obtain a rough estimate of the CAR, any inaccuracies in this estimation may lead to incorrect model selection in the subsequent steps, thereby reducing the accuracy of the final predictions. This motivates us to explore whether it is possible to integrate the strengths of the original model and those trained on sub-datasets into a single model.
@@
To address this, we propose the Multi-Teacher Knowledge Distillation (MTKD) framework, as illustrated in the green box at the bottom of Figure 3. In the O-P strategy, we have already trained the models $\mathcal{M}_{O}$, $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$. Building on this, we further train a student model $\mathcal{M}_{S}$. First, we initialize the student model $\mathcal{M}_{S}$ using the parameters from $\mathcal{M}_{O}$, and then use $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$ as teacher models to perform KD. For each input image pair, we select the appropriate teacher model according to its CAR to supervise the student model. In this framework, the student model $\mathcal{M}_{S}$ is simultaneously supervised by the ground truth labels and the CM information from the teacher models across different CAR partitions.
@@
During the testing phase, only the student model $\mathcal{M}_{S}$ is used for inference, thereby improving the model’s CD performance across different CAR ranges without introducing any additional computational cost.
4.3 Loss Function
@@
For training the original model $\mathcal{M}_{O}$ as well as the teacher models $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$, we employ the standard binary cross-entropy loss:
$\mathcal{L}_{\text{CE}}=\frac{1}{N}\sum_{i=1}^{N}(-Y(i)\log(\hat{Y}(i))-(1-Y(i))\log(1-\hat{Y}(i))),$ (3)
@@
During the training of the student model $\mathcal{M}_{S}$, we dynamically select the appropriate teacher model according to the corresponding CAR range. The knowledge distillation loss is defined as the mean squared error (MSE) between the outputs of the student and the selected teacher at the CM layer:
$\mathcal{L}_{\text{KD}}=\frac{1}{N}\sum_{i=1}^{N}\left(\text{CM}_{S}(i)-\text{CM}_{\mathcal{T}}(i)\right)^{2},\quad\mathcal{T}\in\{T_{S},T_{M},T_{L}\},$ (4)
where $\text{CM}_{S}(i)$ and $\text{CM}_{\mathcal{T}}(i)$ denote the outputs at the CM layer for the student and the selected teacher, respectively. The overall loss for training $\mathcal{M}_{S}$ is a weighted sum of the above objectives:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{KD}},$ (5)
@@
where $\lambda$ is used to balance the cross-entropy loss and the distillation loss.
Table 2: Experimental results on JL1-CD test set.
Method Strategy mIoU mRec mPrec mFscore Method Strategy mIoU mRec mPrec mFscore
- 57.08 61.90 86.40 61.28 - 63.79 69.54 84.77 69.19
O-P 50.05 53.64 95.28 52.00 O-P 60.93 64.72 91.03 65.46
FC-EF MTKD 61.87 68.45 81.90 67.33 FC-Siam-Conc MTKD 64.30 70.58 83.76 69.88
- 61.30 66.03 86.45 66.34 - 66.76 81.71 74.73 74.02
O-P 60.93 64.72 91.03 65.46 O-P 64.56 78.47 72.13 71.25
FC-Siam-Diff MTKD 63.75 69.57 83.95 69.49 STANet MTKD 67.92 82.07 76.24 75.10
- 71.25 78.91 84.53 77.33 - 68.97 74.87 85.06 75.25
O-P 71.06 78.37 84.28 77.21 O-P 71.39 78.60 83.36 77.98
IFN MTKD 72.72 80.28 84.66 78.80 SNUNet MTKD 71.12 78.27 84.96 77.56
- 67.22 74.47 83.71 73.37 - 69.47 75.58 84.46 75.57
O-P 69.41 76.29 84.02 75.77 O-P 68.87 74.74 84.90 74.86
BIT MTKD 68.86 75.49 84.71 74.88 ChangeStar (FarSeg) MTKD 69.87 75.80 85.78 76.03
- 64.85 69.18 88.26 70.19 - 73.51 80.46 86.33 79.70
O-P 64.68 69.05 87.23 70.08 O-P 72.58 79.16 86.33 78.79
ChangeStar (UPerNet) MTKD 65.10 70.26 87.69 70.58 ChangeFormer (MiT-b0) MTKD 73.78 80.07 86.90 79.98
- 73.05 79.70 86.95 79.22 - 71.04 78.77 83.05 77.74
O-P 73.45 79.19 87.46 79.41 O-P 72.22 79.93 83.49 78.76
ChangeFormer (MiT-b1) MTKD 73.92 80.43 86.89 80.18 TinyCD MTKD 72.55 80.98 83.17 79.26
- 63.64 69.77 83.43 69.39 - 74.85 81.84 86.09 80.98
O-P 69.05 76.53 83.05 75.66 O-P 75.29 81.40 87.06 81.32
HANet MTKD 67.67 74.39 84.38 73.92 Changer (MiT-b0) MTKD 75.35 81.76 87.18 81.28
- 75.94 81.99 87.74 81.93 - 68.37 75.15 83.43 74.54
O-P 75.42 81.67 87.13 81.43 O-P 70.76 77.42 83.86 77.01
Changer (MiT-b1) MTKD 76.15 82.85 86.98 82.13 Changer (r18) MTKD 69.45 77.26 81.50 75.86
- 62.31 69.23 80.91 67.83 - 64.58 71.16 83.12 70.13
O-P 71.80 79.76 83.15 78.23 O-P 70.19 77.43 83.99 76.16
Changer (s50) MTKD 62.96 69.65 81.76 68.52 LightCDNet MTKD 65.99 72.44 83.86 71.48
- 73.37 80.31 85.33 79.65 - 73.54 79.54 87.89 79.47
O-P 72.95 79.71 85.50 79.12 O-P 73.61 79.17 88.10 79.45
CGNet MTKD 73.82 80.32 86.33 79.91 BAN MTKD 73.95 80.26 87.12 79.92
- 75.05 80.24 89.82 80.76
O-P 76.69 83.48 87.27 82.52
TTP MTKD 76.85 82.99 88.05 82.56
Figure 4: Visual comparison on the JL1-CD dataset. Each row, from left to right, represents the following: pre-event image, post-event image, ground truth, output from the original model, output from the O-P strategy, and output from the MTKD framework. Red denotes missed detections (FN), while blue indicates false alarms (FP).
Figure 5: mIoU of HANet, ChangeFormer-MiT-b1, and TTP across different CAR ranges. In each plot, the left y-axis represents CAR size, and the right y-axis represents mIoU.
5 Experiment
To comprehensively verify the validity of our JL1-CD dataset and the superiority of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The benchmark methods, implementation details and evaluation metrics used in the experiments are introduced in Appendix B, C, and D , respectively.
5.1 Experimental Results
Quantitative Comparison. Table 2 summarizes the numerical results of mIoU, mRecall (mRec), mPrecision (mPrec), and mF1-score (mFscore) for all methods on the JL1-CD test set, trained under the original, O-P, and MTKD strategies. As shown in the table, the MTKD framework can improve the performance of all the algorithms to a certain extent. However, several algorithms, such as STANet, IFN, and Changer-MiT-b1, perform worse under the O-P strategy than under MTKD, and in some cases, even worse than their original models. This mainly occurs because inaccurate CAR estimation by the original model leads to the selection of an inappropriate teacher model, thereby degrading performance. More qualitative and quantitative analysis of the performance degradation caused by O-P are provided in Appendix F. This finding highlights the potential drawbacks of the O-P strategy in certain scenarios, but it also demonstrates that our MTKD method can enhance student models’ capabilities beyond those of the teacher models trained on partitioned datasets.
Visual Comparison. In remote sensing CD, four main challenges are typically encountered: false alarms, missed detections, internal density, and boundary completeness. To visually assess the effectiveness of the O-P and MTKD strategies in addressing these challenges, we present the visual results of several algorithms in Figure 4, where red indicates missed detections and blue indicates false alarms. Quantitative analysis reveals that both strategies effectively reduce false alarm and missed detection rates while improving internal consistency and boundary precision. Notably, MTKD provides greater improvements over O-P.
Results on Different CAR Partitions. We select one model from each of the CNN, Transformer, and FM architectures and evaluate their performance across different CAR ranges. Figure 5 summarizes the mIoU results of HANet, ChangeFormer-MiT-b1, and TTP on the JL1-CD test set. The images are sorted by CAR in ascending order and divided into five equal partitions. Each bar in the figure represents the lower and upper bounds of CAR for each partition, while the line graph indicates the mIoU across the different partitions. The figure indicates that the detection performance of O-P and MTKD may actually decrease for images with high CAR (e.g., ChangeFormer-MiT-b1 and TTP). However, O-P and MTKD significantly enhance model performance on images with low CAR. In the first partition of the test set (CAR range 0.02% to 0.88%), O-P improves the mIoU for the three algorithms by 8.15%, 3.21%, and 6.55%, respectively, while MTKD improves it by 4.42%, 3.36%, and 2.85%. These results show that by partitioning small CAR regions in the O-P strategy and transferring the knowledge from the corresponding teacher model to the student model in MTKD, our methods achieve substantial improvements in the accuracy of detecting tiny changes.
Comparison of Results on Change and No-Change Classes. We further select some algorithms and compare their detection performance on change and no-change regions. The experimental results are summarized in Table 4. Taking the original models as baselines, we find that MTKD yields consistently larger IoU and Fscore gains on change areas compared to no-change areas, indicating that the MTKD framework is more sensitive to the detection of change areas. For change regions, except CGNet, MTKD notably improves recall more than precision (sometimes even at the cost of reduced precision), suggesting that MTKD is particularly effective at mitigating missed detections. More experimental results on the JL1-CD dataset are provided in Appendix E.
Table 3: Comparison of detection results on change and no-change classes.
Method Class IoU Rec Prec Fscore
unchanged +0.10 -0.60 +0.65 +0.06
SNUNet changed +4.21 +7.38 -0.86 +4.54
unchanged +0.07 +0.12 -0.01 +0.05
ChangeFormer (MiT-b1) changed +1.68 +1.35 -0.11 +1.86
unchanged +0.30 +0.29 +0.08 +0.20
TinyCD changed +2.72 +4.13 +0.16 +2.85
unchanged +0.02 -0.03 -0.04 -0.06
CGNet changed +0.88 +0.03 +2.04 +0.59
unchanged +0.12 +0.23 -0.09 +0.07
BAN changed +0.70 +1.21 -1.46 +0.82
unchanged +0.23 -0.19 +0.45 +0.20
TTP changed +3.36 +5.69 -3.99 +3.39
Table 4: Impact of different numbers of teacher models.
Method Strategy $\mathcal{M}_{T}$ mIoU mFscore
3 75.29 (+0.44) 81.32 (+0.34)
O-P 2 75.44 (+0.59) 81.51 (+0.53)
3 75.35 (+0.50) 81.28 (+0.30)
Changer (MiT-b0) MTKD 2 75.72 (+0.87) 81.66 (+0.68)
3 72.95 (-0.42) 79.12 (-0.53)
O-P 2 73.56 (+0.19) 79.92 (+0.27)
3 73.82 (+0.45) 79.91 (+0.26)
CGNet MTKD 2 73.78 (+0.41) 79.89 (+0.24)
3 76.69 (+1.64) 82.52 (+1.76)
O-P 2 76.65 (+1.60) 82.49 (+1.73)
3 76.85 (+1.80) 82.56 (+1.80)
TTP MTKD 2 76.31 (+1.26) 82.22 (+1.46)
5.2 Robustness of MTKD
Impact of Different Numbers of Teacher Models on MTKD Performance. In the previous subsection, we have demonstrated that the MTKD framework based on three teacher models can effectively improve the performance of various CD models. To alleviate the resource consumption associated with training these models and testing under the O-P strategy, we further explore the performance of the MTKD framework with only two teacher models. Except for the reduced number of teacher models, all other experimental settings remain unchanged. Using the original models as baselines, Table 4 compares the metrics of models trained with the O-P strategy and the MTKD framework under different numbers of teacher models. The results indicate that two teacher models can still achieve notable performance improvements over the baseline. In certain cases, the 2-teacher configuration even outperforms the 3-teacher setup, indicating the flexibility of our approach.
Table 5: Experimental results on SYSU-CD test set.
Method Strategy mIoU mRec mPrec mFscore
- 70.13 79.20 83.41 77.80
O-P 70.25 79.28 83.88 77.81
SNUNet MTKD 70.61 79.85 83.15 78.29
- 71.19 79.29 87.43 78.41
O-P 71.76 79.84 86.62 78.80
ChangeFormer (MiT-b0) MTKD 72.16 80.63 86.63 79.44
- 76.09 84.59 87.08 82.72
O-P 75.97 84.54 86.20 82.66
TTP MTKD 76.11 83.93 87.77 82.77
Performance of MTKD on the SYSU-CD Dataset. We further evaluate the effectiveness of the MTKD framework on the SYSU-CD dataset. CNN-based SNUNet, Transformer-based ChangeFormer-MiT-b0, and FM-based TTP models are selected. As summarized in Table 5, after MTKD optimization, all the models achieve higher mIoU and mFscore values, with the most notable improvement on ChangeFormer-MiT-b0 (mIoU increased by 0.97% and mFscore increased by 1.03%). However, the O-P strategy does not yield better results on TTP compared to its original version. Overall, MTKD demonstrates substantial effectiveness on the SYSU-CD dataset, further validating its robustness. More experimental results on the SYSU-CD dataset are provided in Appendix G.
6 Conclusion
In this work, we present JL1-CD, a new benchmark dataset for CD, which substantially augments existing datasets by providing sub-meter resolution, a diverse array of change types, and a large dataset scale. The MTKD framework is further proposed, which consistently improves the performance of mainstream CD algorithms with various architectures and parameter scales, without incurring additional computational cost during inference. One potential limitation of MTKD is that the number of teacher models and the CAR thresholds are set empirically. Future work includes adaptive selection of the number of teacher models and partition thresholds based on different scenarios to further enhance the flexibility and performance of the MTKD framework.
Appendix
Appendix A CAR Distribution
Figure 6 presents sample images from the JL1-CD dataset with varying values of the Change Area Ratio (CAR). We further provide a quantitative analysis of the CAR distributions for the JL1-CD and SYSU-CD datasets in Figure 7. The CAR spans from 0 to 100% in these datasets, which poses significant challenges for model learning, and it serves as the primary motivation behind our proposed O-P strategy and MTKD framework.
Figure 6: Sample images with different change area ratios (CAR). Each column represents a specific CAR scenario: (a) 0.00%; (b) 19.98%; (c) 39.93%; (d) 59.96%; (e) 80.25%; and (f) 100.00%.
(a) JL1-CD
(b) SYSU-CD
Figure 7: CAR distribution of the JL1-CD and SYSU-CD datasets.
Appendix B Benchmark Methods
To comprehensively verify the validity of our JL1-CD dataset and the effectiveness of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The selected models, along with their corresponding backbones, parameter sizes, and computational complexities, are summarized in Table 6. Based on the backbone architecture, the models are categorized into three groups: Alice Blue for CNN-based models, Light Cyan for Transformer-based models, and Lavender Blue for FM-based models, encompassing almost all mainstream architectures. FLOPs is calculated with an input image of size 512 × 512. As shown in the table, the selected models span a wide range of sizes, from lightweight models with fewer than 1M parameters, such as TinyCD and LightCDNet, to the latest SOTA model TTP with over 360M parameters. This wide range of model sizes also allows us to verify the universality of the O-P and MTKD methods across models with different backbones and scales.
Table 6: Benchmark methods and the corresponding implementation details.